OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 240 and a standard deviation of 32. Convert 144 to a z-score. Round your answer to the nearest hundredth as needed.

OpenStudy (cwrw238):

(144 - 240) / 32 = -96 / 32 = -3